# Representational Stability of Truth in Large Language Models

### v2

@@ 7 Conclusion
This study introduces representational stability as an approach for examining how LLMs internally encode and preserve distinctions between True, False, and Neither statements. By combining controlled label perturbations with a representation-based probe, we show that LLMs display a coherent but uneven geometry of truth. Factual representations are generally well structured, but plausible yet not True statements unseen during training (i.e., unfamiliar Synthetic Neither statements) most readily disrupt that structure. These effects are consistent across architectures and domains, revealing that epistemic familiarity of the LLM with the content, rather than linguistic similarity or model scale, determines the stability of veracity representations. These findings underscore the importance of evaluating not only what LLMs output, but also the reliability of the veracity representations. Understanding and reinforcing this internal coherence offers a path toward models that are not only accurate in response but also epistemically stable in representation.
@@ Acknowledgments
Data availability
True, False, and Synthetic statements are available at https://huggingface.co/datasets/carlomarxx/trilemma-of-truth. Fictional statements are available at https://huggingface.co/datasets/samanthadies/representational_stability. The code used to generate the Noise activations can be found at https://github.com/samanthadies/representational_stability.
Code availability
The code used to generate activations, generate the Noise activations, and run the stability experiments can be found at https://github.com/samanthadies/representational_stability.
@@ Appendix A Notation
Symbol Description
$\mathcal{M}$ A given LLM.
NN Number of statements in the dataset.
$\mathcal{S}=\{s_{i}\}_{i=1}^{N}$ Set of natural-language statements.
$y_{i}$ Ground-truth label of $s_{i}$: $\{\texttt{True},\,\texttt{False},\,\texttt{Neither}\}$.
$l$ Index of the LLM layer used for activation extraction.
$\mathbf{z}_{i}^{(l)}$ Activation vector for statement $s_{i}$ at layer $l$.
$\mathcal{D}=\{(\mathbf{z}_{i}^{(l)},y_{i})\}_{i=1}^{N}$ Full activation–label dataset.
$\mathcal{D}_{\mathrm{train}},\ \mathcal{D}_{\mathrm{cal}},\ \mathcal{D}_{\mathrm{test}}$ Train, calibration, and test splits of $\mathcal{D}$.
$y^{\prime}_{i}\in\{+1,-1\}$ Baseline binary label: +1+1 for True, and 1-1 for False/Neither.
$f(\mathbf{z})=\vec{w}\cdot\mathbf{z}+b$ Linear decision function of the baseline probe.
$\vec{w}$ Learned “truth direction” for the baseline probe.
$b$ Bias term of the baseline probe.
$\hat{y}_{i}$ Predicted label of the baseline probe for $s_{i}$.
$\mathcal{B}_{\mathrm{true}}=\{s_{i}\mid\mathbf{z}_{i}^{(l)}\in\mathcal{D}_{\mathrm{test}},\ y_{i}=\mathtt{True},\ \hat{y}_{i}=+1\}$ True statements classified as true (baseline belief set).
$\mathcal{N}$ Set of all Neither statements.
$\mathcal{N}_{1},\ \mathcal{N}_{0}$ Partition of $\mathcal{N}$: those re-labeled as True ($\mathcal{N}_{1}$) versus Not True ($\mathcal{N}_{0}$).
$y^{\prime\prime}_{i}\in\{+1,-1\}$ Perturbed binary label under a given $\mathcal{N}_{1}/\mathcal{N}_{0}$ split.
$\vec{w}^{\prime},\,b^{\prime}$ Truth direction and bias learned by the perturbed probe.
$\hat{y}^{\prime}_{i}$ Predicted label of the perturbed probe.
$\mathcal{B}^{\prime}_{\mathrm{true}}=\{s_{i}\mid\mathbf{z}_{i}^{(l)}\in\mathcal{D}_{\mathrm{test}},\ y_{i}=\mathtt{True},\ \hat{y}^{\prime}_{i}=+1\}$ True statements classified as true under the perturbed probe.
@@ B.1 Data Generation
We use statements from the City Locations, Medical Indications, and Word Definitions datasets introduced in [savcisens2025trilemma]. City statements take the form “The city of [city] is (not) located in [country],” (omitting “The city of” when redundant). Medical statements follow “[drug] is (not) indicated for the treatment of [disease/condition].” Word Definition statements draw from three templates: “[word] is (not) a [instanceOf],” “[word] is (not) a type of [typeOf],” and “[word] is (not) a synonym of [synonym].”
@@ B.1.1 True, False, and Synthetic Statements
We take the True, False, and Synthetic statements from the datasets introduced in [savcisens2025trilemma]. All statements are constructed with both affirmative and negated forms. Synthetic entities are generated using a Markov-chain–based name generator (namemaker [^4] https://github.com/Rickmsd/namemaker) and undergo multi-stage filtering, including database checks, model tagging, and web-search validation, to ensure no accidental overlap with real entities. Validated names are then paired to form grammatically coherent but semantically meaningless statements that follow each template. Because Synthetic entities do not exist and cannot have appeared in training corpora, LLMs have no basis for assigning them a truth value. Accordingly, these statements function as Neither cases: unknown claims for which belief should be suspended rather than confidently classified as true or false.
@@ B.1.2 Fictional Statements
In addition to Synthetic statements, which represent unseen and unknown claims, we construct new sets of Fictional statements for all three domains. Fictional statements also function as Neither statements in our experiments as they reference entities that do not exist in the real world and therefore lack real-world truth value. However, unlike Synthetic statements, many Fictional entities are likely to have appeared in LLM training corpora. [^5] For later analyses, we additionally annotate fictional statements with their within-universe factual status (Fictional True or Fictional False), but this labeling is not used in the primary True vs. Not True classification tasks. As such, they represent a complementary form of Neither: claims that an LLM may recognize, but that still lie outside the true–false axis relevant to factual grounding.
@@
Fictional drug and disease statements are drawn from (1) NeoEncyclopedia Wiki [fandom_fictional_diseases, fandom_fictional_toxins]; (2) ChemEurope’s List of Fictional Medicines and Drugs [chemeurope_fictional_medicines_drugs]; and (3) The Thackery T. Lambshead Pocket Guide to Eccentric & Discredited Diseases [tomasula2004lambshead]. Drug–disease pairs are included when a treatment relationship exists within the fictional source.
@@
Fictional lexical items are compiled from (1) Gobblefunk [beelinguappDahlDictionary]; (2) Dothraki [conlangDothrakiInitial]; and (3) Na’vi [dict_navi_online_dictionary]. Dothraki and Na’vi have formal linguistic structure, whereas Gobblefunk is a playful neologistic extension of English.
@@ B.1.3 Noise
The Noise statements contains no linguistic content. We generate $n_{\mathrm{noise}}=0.10\cdot|\mathcal{D}|$ random activation sequences by sampling from a multivariate Gaussian with per-feature mean, standard deviation, and sequence-length distribution matched to the LLM activations. These distributionally consistent but non-semantic sequences serve as a control, allowing us to test whether observed representational differences arise from semantic content or from statistical variation in activation space.
@@ B.2 Data Splits for Probing Experiments
Definitions $6488$ $(0.54)$ $2514$ $(0.21)$ $3041$ $(0.25)$ $12043$ $(1.00)$
Table A2 summarizes the dataset partitions used for all probing experiments. Each dataset is split exclusively into training, calibration, and test sets to prevent data leakage. Approximately $55\%$ of statements are used for training, $20\%$ for calibration, and $25\%$ for testing. We use identical splits across all probes and all LLMs, enabling direct comparison of representational stability under matched data conditions.
@@ Appendix C LLMs
Table A3 lists the sixteen open-source LLMs used in our stability experiments. The set spans four major model families, Gemma, Llama, Mistral, and Qwen, with between about $3$ billion to about $15$ billion parameters and release dates between February and September $2024$. For each family, we include both base (pre-trained) and chat-tuned variants to capture differences introduced by instruction fine-tuning. Together, these models provide a representative cross-section of current decoder-only architectures varying in scale, origin, and training objectives.
Table A3: LLMs used in the stability experiments. We list the official names of the LLMs according to the HuggingFace repository [wolf2020transformers]. We further specify the shortened name we use to refer to each of the models, whether it is the base, pre-trained model or a chat-tuned version, the nubmer of decoders, the number of parameters, the release date, and the source of the model. Finally, we report the layers with the best separation between True and Not True statements for the City Locations (“C”), Medical Indications (“M”), and Word Definitions (“W”) datasets. The LLMs are publicly available through HuggingFace [wolf2020transformers].
@@
Llama-33-88b llama-3-8b Base $32$ $8.03$ B C: $18$, M: $17$, W: $17$ Jul $23$, $2024$ Meta
@@
Appendix D Representations of True, False, and Neither by LLM
Supplementary Figures A1-A16 show the pairwise activation distance matrices for all sixteen LLMs. Three general representational patterns emerge. The first, observed in _gemma-2-9b (Fig. A1) and gemma-2-9b (Fig. A10), shows Fictional and Synthetic statements clustering near True and False statements, with Noise forming a distinct outlier. The second, present in _gemma-7b (Fig. A2), gemma-7b (Fig. A10), _qwen-2.5-14b (Fig. A8), qwen-2.5-14b (Fig. A15), and _qwen-2.5-7b (Fig. A9), exhibits Synthetic statements close to True and False, Fictional statements clearly separated, and Noise positioned slightly closer to the True/False/Synthetic cluster. The third, seen in the remaining nine models, features Synthetic statements aligned with True and False, while both Fictional and Noise statements occupy distinct and distant regions. Except for _qwen-2.5-7b (which follows the second pattern; Fig. A9) and qwen-2.5-7b (the third; Fig. A16), base and chat versions of each model display qualitatively similar representational structures.
@@ Appendix D Representations of True, False, and Neither by LLM
Figure A15: Wasserstein distance between activations for qwen-2.5-14b. Pairwise Wasserstein distances between activation distributions of True, False, Synthetic, Fictional, and Noise statements for the (a) City Locations, (b) Medical Indications, and (c) Word Definitions datasets. Synthetic statements are represented similarly to True and False statements, while Fictional statements are represented distinctly from all other statements.
@@
Appendix E Exploring Additional Probes

### v3

@@ 7 Conclusion
We show that epistemic familiarity is a key determinant of belief stability under semantic reframing. By linking internal representations and behavioral responses under matched perturbations, P-StaT reveals that unfamiliar Synthetic content consistently induces the largest epistemic retractions across LLMs and domains. These results demonstrate that evaluating belief stability under controlled perturbations of Neither content provides a principled complement to accuracy-based factuality metrics and a step toward more epistemically reliable language models.
8 Limitations
Our perturbations focus on a specific notion of epistemic familiarity and a limited set of factual domains. While this design enables controlled comparisons across representational and behavioral settings, it does not exhaustively cover all types of semantic variation. Extending P-StaT to other forms of epistemic ambiguity, such as disputed claims, probabilistic beliefs, or evolving facts, would further test its generality. In addition, our analysis considers fixed LLM parameters. Although P-StaT isolates how semantic perturbations interact with existing internal representations, it does not address how belief stability may change in settings where representations themselves evolve over time. Finally, retraction rates depend on accuracy on True statements, so our notion of stability reflects both robustness to perturbation and an LLM’s propensity to assert truth in the baseline case. While we emphasize epistemic retractions as a primary signal of instability, other applications may require alternative notions of stability.
9 Ethical Considerations
Our study examines the conditions under which perturbations can systematically destabilize what an LLM considers true. For example, we find that unfamiliar synthetic content induces the largest epistemic retractions, producing up to 32.732.7% retractions in representational evaluations and up to 36.336.3% in behavioral evaluations. This could inform efforts to undermine LLM reliability and has implications for trust in LLMs.
While our framework could be misused to deliberately destabilize LLM beliefs, our intent is diagnostic rather than adversarial: P-StaT is designed to identify epistemic vulnerabilities in order to inform more robust evaluation and LLM design. We do not propose interventions or belief manipulation techniques, and all experiments are conducted on fixed, open-source LLMs in offline settings. More broadly, our findings highlight a structural limitation of current LLMs rather than a prescription for exploiting it. P-StaT is intended solely for research and diagnostic evaluation of pretrained LLMs and is not designed for deployment, belief steering, or real-world decision-making systems.
9.1 Data availability
We use the True, False, and Synthetic statements available at https://huggingface.co/datasets/carlomarxx/trilemma-of-truth under a CC-BY-4.0 license. Fictional statements are available at https://huggingface.co/datasets/samanthadies/representational_stability. The code used to generate the Noise activations can be found at https://github.com/samanthadies/representational_stability.
9.2 Code availability
We release the code used to generate activations, generate the Noise activations, and run the P-StaT stability experiments under an MIT License at https://github.com/samanthadies/representational_stability. ChatGPT and Copilot were used to help write initial versions of experiment and plotting scripts, and helped clean and comment the code. All AI-generated content was reviewed and verified by the authors.
All models used in this work are publicly available for research use under their respective licenses (MIT License for sAwMIL; Gemma for Gemma-7b, Gemma-7b-it, Gemma-2-9b, and Gemma-2-9b-it; llama3.1 for Llama-3.1-8b and Llama-3.1-8b-Instruct; llama3.2 for Llama-3.2-3b, Llama-3.2-3b-Instruct; llama3 for Llama3-Med42-8b; Bio-Medical-Llama-3-8b LLM License for Bio-Medical-Llama-3-8b; apache-2.0 for Mistral-7B-v0.3, Mistral-7B-Instruct-v0.3, Qwen2.5-7B, Qwen2.5-7B-Instruct, Qwen2.5-14B, and Qwen2.5-14B-Instruct). Marks and Tegmark did not specify a license for the Mean Difference probe [20].
@@ Appendix A Notation
Symbol Description
$\mathcal{M}$ A fixed large language model (LLM).
$l$ Layer index used for activation extraction.
$\mathcal{S}=\{s_{i}\}_{i=1}^{N}$ Set of $N$ natural-language statements.
$y_{i}$ Ground-truth veracity label of $s_{i}$, $y_{i}\in\{\texttt{True},\texttt{False},\texttt{Neither}\}$.
$\mathcal{N}$ Set of all Neither statements: $\mathcal{N}=\{s_{i}\mid y_{i}=\texttt{Neither}\}$.
$\mathcal{N}_{\mathrm{fam}},\ \mathcal{N}_{\mathrm{unf}}$ Partition of $\mathcal{N}$ into epistemically familiar (Fictional) and unfamiliar (Synthetic) subsets.
$\mathcal{N}_{\mathrm{fam}}^{(T)},\ \mathcal{N}_{\mathrm{fam}}^{(F)}$ Canonically true and false subsets of familiar fictional statements (used in Fictional(T)).
$\phi_{\mathcal{M},l}$ Representation map from a statement to its layer-$l$ hidden representations under $\mathcal{M}$.
$\mathbf{z}_{i}^{(l)}$ Layer-$l$ representation of statement $s_{i}$, $\mathbf{z}_{i}^{(l)}=\phi_{\mathcal{M},l}(s_{i})$.
$\mathcal{D}$ Dataset of statements, representations, and labels: $\mathcal{D}=\{(s_{i},\mathbf{z}_{i}^{(l)},y_{i})\}_{i=1}^{N}$.
$\mathcal{D}_{\mathrm{train}},\ \mathcal{D}_{\mathrm{test}}$ Training and test splits of $\mathcal{D}$.
$\Theta$ Semantic interpretation specifying which Neither statements are treated as compatible with truth.
$\mathcal{N}_{\Theta}\subseteq\mathcal{N}$ Subset of Neither statements treated as compatible with truth under $\Theta$.
$g(\cdot,\Theta)$ Evaluation function mapping a statement and interpretation to a binary judgment: $g(s,\Theta)\in\{\texttt{True},\texttt{Not True}\}$.
$g_{\mathrm{p}}$ Representational instantiation of $g$ via a probe over $\phi_{\mathcal{M},l}(s)$.
$g_{\mathrm{zs}}$ Behavioral instantiation of $g$ via zero-shot prompting with belief context $C_{\Theta}$.
$C_{\Theta}$ The belief context inserted into a zero-shot prompt to instantiate perturbation $\Theta$.
$\Theta_{0}$ Baseline semantic interpretation (no Neither treated as true).
$\mathcal{B}^{\Theta_{0}}_{\mathrm{true}}$ Baseline belief set: $\{s_{i}\mid y_{i}=\texttt{True},\ g(s_{i},\Theta_{0})=\texttt{True}\}$.
$\mathcal{B}^{\Theta}_{\mathrm{true}}$ Belief set under perturbed interpretation $\Theta$.
$\mathcal{R}$ Epistemic retractions: $\mathcal{R}=\mathcal{B}^{\Theta_{0}}_{\mathrm{true}}\setminus\mathcal{B}^{\Theta}_{\mathrm{true}}$.
$\mathcal{E}$ Epistemic expansions: $\mathcal{E}=\mathcal{B}^{\Theta}_{\mathrm{true}}\setminus\mathcal{B}^{\Theta_{0}}_{\mathrm{true}}$.
@@ B.1 Data Generation
We use statements from the City Locations, Medical Indications, and Word Definitions datasets introduced in [23]. City statements take the form “The city of [city] is (not) located in [country],” (omitting “The city of” when redundant). Medical statements follow “[drug] is (not) indicated for the treatment of [disease/condition].” Word Definition statements draw from three templates: “[word] is (not) a [instanceOf],” “[word] is (not) a type of [typeOf],” and “[word] is (not) a synonym of [synonym].” No personal data, identifying information, or user-generated content is included.
@@ B.1.1 True, False, and Synthetic Statements
We take the True, False, and Synthetic statements from the datasets introduced in [23]. All statements are constructed with both affirmative and negated forms. Synthetic entities are generated using a Markov-chain–based name generator (namemaker [^5] https://github.com/Rickmsd/namemaker.) and undergo multi-stage filtering, including database checks, model tagging, and web-search validation, to ensure no accidental overlap with real entities. Validated names are then paired to form grammatically coherent but semantically meaningless statements that follow each template. Because Synthetic entities do not exist and cannot have appeared in training corpora, LLMs have no basis for assigning them a truth value. Accordingly, these statements function as Neither cases: unknown claims for which belief should be suspended rather than confidently classified as true or false.
@@ B.1.2 Fictional Statements
In addition to Synthetic statements, which represent unseen and unknown claims, we construct new sets of Fictional statements for all three domains. Fictional statements also function as Neither statements in our experiments as they reference entities that do not exist in the real world and therefore lack real-world truth value. However, unlike Synthetic statements, many Fictional entities are likely to have appeared in LLM training corpora. [^6] For later analyses, we additionally annotate fictional statements with their within-universe factual status (Fictional (T) or Fictional (F)), but this labeling is not used in the primary True vs. Not True classification tasks. As such, they represent a complementary form of Neither: claims that an LLM may recognize, but that still lie outside the true–false axis relevant to factual grounding.
@@
Fictional drug and disease statements are drawn from (1) NeoEncyclopedia Wiki [10, 11]; (2) ChemEurope’s List of Fictional Medicines and Drugs [6]; and (3) The Thackery T. Lambshead Pocket Guide to Eccentric & Discredited Diseases [28]. Drug–disease pairs are included when a treatment relationship exists according to the fictional source.
@@
Fictional lexical items are compiled from (1) Gobblefunk [3]; (2) Dothraki [24]; and (3) Na’vi [8]. Dothraki and Na’vi have formal linguistic structure, whereas Gobblefunk is a playful neologistic extension of English.
@@ B.1.3 Noise
The Noise statements contain no linguistic content. We generate $n_{\mathrm{noise}}=0.10\cdot|\mathcal{D}|$ random activation sequences by sampling from a multivariate Gaussian with per-feature mean, standard deviation, and sequence-length distribution matched to the LLM activations. These distributionally consistent but non-semantic sequences serve as a control, allowing us to test whether observed representational differences arise from semantic content or from statistical variation in activation space.
@@ B.2 Data Splits for Probing Experiments
Word Definitions $6488$ $(0.54)$ $2514$ $(0.21)$ $3041$ $(0.25)$ $12043$ $(1.00)$
Table A2 summarizes the partitions used for all experiments. Each dataset is split exclusively into training, calibration, and test sets to prevent data leakage. Approximately $55\%$ of statements are used for training, $20\%$ for calibration, and $25\%$ for testing. We use identical splits in all conditions.
@@ Appendix C LLMs
Table A3 lists the sixteen open-source LLMs used in our experiments. The set spans four major model families, Gemma, Llama, Mistral, and Qwen, with approximately $3$ billion to $15$ billion parameters and release dates between February and September $2024$. For each family, we include both base (pre-trained) and chat-tuned variants. Together, these LLMs provide a representative cross-section of current decoder-only architectures varying in scale, origin, and training objectives.
@@
Llama-33-88b llama-3.1-8b Base $32$ $8.03$ B C: $18$, M: $17$, W: $17$ Jul $23$, $2024$ Meta
@@
Table A3: LLMs used in the stability experiments. We list the official names of the LLMs according to the HuggingFace repository [34]. We further specify the shortened name we use to refer to each of the LLMs, whether it is the base, pre-trained LLM or a chat-tuned version, the number of decoders, the number of parameters, the release date, and the source of the LLM. Finally, we report the layers with the best separation between True and Not True statements for the City Locations (“C”), Medical Indications (“M”), and Word Definitions (“W”) datasets. The LLMs are publicly available through HuggingFace [34].
Appendix D LLM-level Results
The main text reports LLM-agnostic stability patterns aggregated across sixteen LLMs. Here we provide the corresponding LLM-level results.
D.1 Linguistic and Representational Structure of Neither Statements (by LLM)
Figures A1-A16 show the pairwise activation distance matrices for all sixteen LLMs. Three general representational patterns emerge. The first, observed in _gemma-2-9b (Fig. A1) and gemma-2-9b (Fig. A10), shows Fictional and Synthetic statements clustering near True and False statements, with Noise forming a distinct outlier. The second, present in _gemma-7b (Fig. A2), gemma-7b (Fig. A11), _qwen-2.5-14b (Fig. A8), qwen-2.5-14b (Fig. A15), and _qwen-2.5-7b (Fig. A9), exhibits Synthetic statements close to True and False, Fictional statements clearly separated, and Noise positioned slightly closer to the True/False/Synthetic cluster. The third, seen in the remaining nine LLMs, features Synthetic statements aligned with True and False, while both Fictional and Noise statements occupy distinct and distant regions. Except for _qwen-2.5-7b (which follows the second pattern; Fig. A9) and qwen-2.5-7b (the third; Fig. A16), base and chat versions are qualitatively similar.
@@ D.1 Linguistic and Representational Structure of Neither Statements (by LLM)
Figure A15: Wasserstein distance between activations for qwen-2.5-14b. Pairwise Wasserstein distances between activation distributions of True, False, Synthetic, Fictional, and Noise statements for the (a) City Locations, (b) Medical Indications, and (c) Word Definitions. Synthetic statements are represented similarly to True and False statements, while Fictional statements are represented distinctly from all other statements.
@@
D.2 Representational Stability under Probing Perturbations (by LLM)
Figure A17 shows how the sAwMIL decision boundaries change under each perturbation, measured by cosine similarity to the baseline True vs. Not True direction and by the associated bias shift. Consistent with the aggregate flip rates (Table 3), Synthetic produces the largest boundary changes across domains, while Fictional, Fictional(T), and Noise yield smaller deviations.
Different LLM families exhibit different degrees of susceptibility to these perturbations. Chat-tuned variants (denoted by leading underscores) tend to exhibit somewhat larger rotations and bias shifts than their base models. An exception is gemma-7b, which shows unusually large shifts in the Word Definitions domain.
Figures A18A20 break down epistemic retractions $\mathcal{R}$ and expansions $\mathcal{E}$ by LLM. Chat-tuned LLMs tend to produce more expansions, while Base models tend to exhibit more retractions. However, these tendencies do not hold uniformly, and overall differences across LLM families are smaller than differences across perturbation types.
Figure A17: Changes in the probe decision boundary under perturbations. Cosine similarity (left column) and bias difference (right column) between the baseline True vs. Not True probe and probes retrained under label perturbations for the (a,b) City Locations, (c,d) Medical Indications, and (e,f) Word Definitions datasets. Each heatmap shows results for sixteen LLMs (columns) and five perturbation conditions (rows). LLMs with leading underscores are Chat models, while those without are Base models. Higher cosine similarity indicates smaller rotations of the learned decision boundary, while bias difference reflects shifts in intercept. Across datasets, probes retrained with the Synthetic perturbation show the largest deviation from the original, particularly in cosine similarity.
Figure A18: Stability of probe predictions under belief context perturbations for City Locations data. Bar plots show for each LLM (x-axis) how often the probe induces epistemic expansions and retractions when retrained under four perturbations: (a) Synthetic, (b) Fictional, (c) Fictional(T), and (d) Noise. Green bars indicate retractions (True to Not True), while purple bars indicate expansions (Not True to True). The left y-axis reports the number of statements with flipped predictions, and the right y-axis reports the corresponding proportions. The Base models exhibit more retractions than the Chat models, and the Chat models induce more expansions.
Figure A19: Stability of probe predictions under belief context perturbations for Medical Indications data. Bar plots show for each LLM (x-axis) how often the probe induces epistemic expansions and retractions when retrained under four perturbations: (a) Synthetic, (b) Fictional, (c) Fictional(T), and (d) Noise. Green bars indicate retractions (True to Not True), while purple bars indicate expansions (Not True to True). The left y-axis reports the number of statements with flipped predictions, and the right y-axis reports the corresponding proportions. The Synthetic perturbation leads to the most instability, and the Fictional and Fictional (T) perturbations result in almost no flips.
Figure A20: Stability of probe predictions under belief context perturbations for Word Definitions data. Bar plots show for each LLM (x-axis) how often the probe induces epistemic expansions and retractions when retrained under four perturbations: (a) Synthetic, (b) Fictional, (c) Fictional(T), and (d) Noise. Green bars indicate retractions (True to Not True), while purple bars indicate expansions (Not True to True). The left y-axis reports the number of statements with flipped predictions, and the right y-axis reports the corresponding proportions. The Synthetic perturbation leads to the most instability, with some LLMs retracting over $50\%$ of their originally True statements.
D.3 Behavioral Stability under Zero-Shot Perturbations (by LLM)
Figures A21A23 break down epistemic retractions $\mathcal{R}$ and expansions $\mathcal{E}$ by LLM. Opposite to the behavior seen in the representational instantiation, Base LLMs tend to produce more expansions, while Chat-tuned variants tend to exhibit more retractions.
Figure A21: Stability of zero-shot beliefs under belief context perturbations for City Locations data. Bar plots show for each LLM (x-axis) how often the zero-shot instantiation induced epistemic expansions and retractions when retrained under three perturbations: (a) Synthetic, (b) Fictional, and (c) Fictional(T). Green bars indicate retractions (True to Not True), while purple bars indicate expansions (Not True to True). The left y-axis reports the number of statements with flipped predictions, and the right y-axis reports the corresponding proportions. The Chat models exhibit more retractions than the Base models.
Figure A22: Stability of zero-shot beliefs under belief context perturbations for Medical Indications data. Bar plots show for each LLM (x-axis), how often the zero-shot instantiation induced epistemic expansions and retractions when retrained under three perturbations: (a) Synthetic, (b) Fictional, and (c) Fictional(T). Green bars indicate retractions (True to Not True), while purple bars indicate expansions (Not True to True). The left y-axis reports the number of statements with flipped predictions, and the right y-axis reports the corresponding proportions. The Chat models exhibit more retractions, while the Base models exhibit more expansions.
Figure A23: Stability of zero-shot beliefs under belief context perturbations for Word Definitions data. Bar plots show for each LLM (x-axis), how often the zero-shot instantiation induced epistemic expansions and retractions when retrained under three perturbations: (a) Synthetic, (b) Fictional, and (c) Fictional(T). Green bars indicate retractions (True to Not True), while purple bars indicate expansions (Not True to True). The left y-axis reports the number of statements with flipped predictions, and the right y-axis reports the corresponding proportions. The Chat models exhibit more retractions, while the Base models exhibit more expansions.
Appendix E Exploring the Mean Difference Probe